# Learning About Structural Errors in Models of Complex Dynamical Systems

### v1

@@ 1 Introduction
Numerical simulation is at the heart of modeling, predicting, and understanding dynamical systems that are too complex to be amenable to analytical solution. Complex dynamical systems here extend from molecular dynamics with quantum effects to the planetary scales of weather and climate. The range of dynamically important scales in these systems can be vast, for example, in case of the atmosphere, extending over 13 orders of magnitude from the micrometers of cloud droplets and aerosols to the tens of thousands kilometers of planetary waves. The number of degrees of freedom that would need to be resolved for a faithful simulation of such systems (e.g., $\gtrsim 10^{21}$ for a typical atmospheric boundary layer flow) often exceeds what will be computationally feasible for the foreseeable future Schneider et al. (2017).
@@
where $f$ may depend on derivatives of the state $X$; hence, the system may represent partial differential equations. The system depends on empirical parameters $\theta_{\textrm{P}}$ that appear in closure models. For example, in large-eddy simulations of turbulent flows, the most energetic “large eddies” are explicitly resolved in the dynamics represented by $f$. The effect of the unresolved scales is modeled by subgrid-scale models, such as the classical Smagorinsky model Smagorinsky (1963), which depend on empirical parameters $\theta_{\textrm{P}}$ (e.g., the Smagorinsky coefficient). Similar semi-empirical models are used in many other fields. They encode domain-specific knowledge, and their parameters $\theta_{\textrm{P}}$ need to be calibrated with data.
@@
The challenge is that simulated observables $\hat{y}$ generally are biased estimates of actual data $y$. The actual data $y$ are affected by measurement error, and the simulated data $\hat{y}$ are affected by structural errors in the approximate dynamical system (1); both $y$ and $\hat{y}$ can also be affected by sampling error. For example, while a general feature of turbulence is to enhance mixing of conserved quantities, turbulent mixing is not always diffusive in character. Therefore, diffusive subgrid-scale models such as the Smagorinsky model are not always structurally correct, especially in convective situations with coherent flow structures Holtslag and Moeng (1991). This can lead to biases that, for example, adversely affect the calibration of model parameters $\theta_{\textrm{P}}$.
The purpose of this paper is to summarize principles of, and algorithms for, learning about structural error models that correct semi-empirical closure models. Wholesale replacement of semi-empirical closure models with neural networks and other deep learning approaches promises to overcome the structural strictures of existing closure models through more expressive models; it has recently received much attention Ling et al. (2016); Weatheritt and Sandberg (2016); Rasp et al. (2018); Maulik et al. (2019); Duraisamy et al. (2019); Zhao et al. (2020); Schmelzer et al. (2020); Brunton et al. (2020). However, existing semi-empirical closure models encode valuable domain-specific knowledge. Learning flexible corrections to these models is often less data hungry, more interpretable, and potentially more generalizable than replacing them wholesale.
@@
Our approach to constructing error models builds upon but goes beyond the classical work of Kennedy and O’Hagan Kennedy and O’Hagan (2001), who accounted for model error through an external bias correction term $\delta(X;\theta_{\textrm{E}})$, parameterized by parameters $\theta_{\textrm{E}}$ and inserted at the boundary between output from a computer model $\hat{y}=\mathcal{G}(\theta_{\textrm{P}})$ and data $y$:
@@
Here, $\mathcal{G}(\theta_{\textrm{P}})=\mathcal{H}[X(\theta_{\textrm{P}})]$ corresponds to solving (1) for the time series of the state $X$, which depends parametrically on $\theta_{\textrm{P}}$, and then applying the observation operator (2); hence, $\mathcal{G}$ is a mapping from the space of model parameters $\theta_{\textrm{P}}$ to the space of observations $y$. The noise $\eta$ represents additional (e.g., observation) errors, assumed to have zero mean. In this approach, the model parameters $\theta_{\textrm{P}}$ remain fixed (i.e., a property of $\mathcal{G}$) while parameters $\theta_{\textrm{E}}$ in the error model $\delta$ are tuned such that the residual $y-\hat{y}$ has a small magnitude and zero mean. This approach of externalizing model error for bias correction has been applied and further expanded in many subsequent papers (e.g., Higdon et al., 2004; Wang et al., 2009; Brynjarsdòttir and O’Hagan, 2014; Trehan et al., 2017; Calvetti et al., 2018). A key advantage of the external model error approach is that the model producing $\hat{y}$ can be treated as a black-box, which facilitates use of this approach across different domains. State variables $X$ and residuals $y-\hat{y}$ form input-output pairs from which the error model $\delta(X;\theta_{\textrm{E}})$ can be learned, for example, with supervised learning [^3] Supervised learning refers to regression or interpolation of data. approaches, usually in an offline setting separate from learning about the model parameters $\theta_{\textrm{P}}$.
@@
    It is difficult to incorporate physical (or other process-based) knowledge or constraints (e.g., conservation laws) in the error model $\delta(X;\theta_{\textrm{E}})$ Brynjarsdòttir and O’Hagan (2014).
@@
To address these drawbacks, a few researchers have started to explore an approach that internalizes model error Strong and Oakley (2014); He and Xiu (2016); Sargsyan et al. (2019). Such an internal model error approach embeds $L$ error models $\delta_{l}(X;\ \theta_{\textrm{I}}^{(l)})$ ($l=1,\dots,L$) within the dynamical system, at the places (e.g., in closure models) where the errors are actually made. Let their collection be written as
@@
Such an approach has found applications, for example, in turbulence modeling Emory et al. (2011); Oliver and Moser (2011); Cheung et al. (2011); Xiao et al. (2016); Pernot and Cailliez (2017); Wang et al. (2017); Wu et al. (2018). By incorporating the structural error models $\delta(X;\ \theta_{\textrm{I}})$ inside the dynamical system, the error models can in principle lead to improved predictions even of quantities that were not used to train the error models. The error models can be learned alongside the model parameters $\theta_{\textrm{P}}$ in an online setting. They also are more amenable to interpretation because they are included in the places where errors are actually made. A potential downside of internalizing model error is that the effects of the structural errors $\delta$ map onto data $y$ only through the resolved state $X$ of the dynamical system, and residuals $y-\hat{y}$ are generally not directly informative about how the structural errors $\delta(X;\ \theta_{\textrm{I}})$ depend on state variables $X$; thus, learning about structural errors $\delta(X;\ \theta_{\textrm{I}})$ can generally not be accomplished with direct supervised learning approaches. Instead, residuals $y-\hat{y}$ only provide indirect information about structural errors $\delta(X;\ \theta_{\textrm{I}})$. Additionally, if derivatives of the dynamical system $f$ with respect to parameters are not easily available, or if the dynamical system is not differentiable, gradient-based methods for learning about the model errors $\delta(X;\ \theta_{\textrm{I}})$ are difficult or impossible to use.
@@
    Gaussian processes, as in Kennedy and O’Hagan Kennedy and O’Hagan (2001);
@@
    Models assembled from dictionaries of terms (e.g., involving differential operators), as in the data-driven discovery of partial differential equations Schaeffer et al. (2013); Rudy et al. (2017); Schaeffer (2017); Zanna and Bolton (2020); de Silva et al. (2020); Lemos et al. (2022);
@@
    Neural networks, for their expressivity Ling et al. (2016); Brenowitz and Bretherton (2018); Rasp et al. (2018); Maulik et al. (2019); Zanna and Bolton (2021); Bretherton et al. (2022);
@@
    Stochastic models, because without a clear scale separation between resolved and unresolved degrees of freedom, theoretical analysis suggests that closure models generally should be stochastic (e.g., Zwanzig, 2001; Lucarini et al., 2014; Franzke et al., 2015; Palmer, 2019);
@@
    Non-local models, because structural errors may be non-local in space Zhou et al. (2021); Zhuang et al. (2021); Du et al. (2012); D’Elia and Gunzburger (2016); Pang et al. (2020), in time Ma et al. (2018); Wang et al. (2020); Lin and Lu (2021), or in both Charalampopoulos and Sapsis (2021).
We will discuss how to learn about such error models both from direct and indirect data. Supervised learning of various types of error models $\delta_{l}$ from direct data has been performed in a number of settings, for example, to discover terms in differential equations or neural network closure models from residual time tendencies that give direct information about the error model that is sought (e.g., Brunton et al., 2016; Wang et al., 2017; Wu et al., 2018; Levine and Stuart, 2022). However, data directly informative about error models, such as high-resolution time tendencies, are not always available. When training an error model on time tendencies, it can also be difficult to ensure both stability of the dynamical system including the error model and satisfaction of physical constraints (e.g., energy conservation) (e.g., Bretherton et al., 2022). Training an error model with indirect data, in an inverse problem rather than supervised learning setting (e.g., Xiao et al., 2016; Schneider et al., 2021; Zhang et al., 2022), can be advantageous in those circumstances. We demonstrate how it can be accomplished.
@@
The principles and algorithms we will discuss are broad purpose and applicable across a range of domains and model complexities. To illustrate their properties in a relatively simple setting, we use two versions of the Lorenz 96 Lorenz (1996) dynamical system: the basic version of the model and its multiscale generalization Fatkullin and Vanden-Eijnden (2004). Section 2 discusses the calibration of internal error models with direct or indirect data and the enforcement of constraints such as conservation properties. Section 3 introduces various ways of constructing error models. Section 4 introduces two Lorenz 96 systems and then proceeds to present various concepts and methods through numerical results for these systems. The Lorenz 96 systems serve as illustrative examples of ergodic systems, for which we avoid trajectory matching and can demonstrate the use of time-averaged statistics, which approximate the expectation with respect to the invariant measure, as data from which to learn about the underlying unknown parameters. Section 5 is organized similarly to the previous section, but concerns non-ergodic systems for which trajectory matching is needed to learn about the underlying unknown parameters; in this context we study a model of the human glucose-insulin system. In Section 6, we state our conclusions.
@@ 2.1.1 Direct Data
as an example. A fine temporal resolution of $X(t)$ is usually needed to approximate $\dot{X}-f(X;\theta_{\textrm{P}})$ and obtain estimates of the error $\delta(X;\theta_{\textrm{I}})$ as a residual. With this method, it becomes challenging to obtain reliable direct data when the trajectories $\dot{X}$ are noisy, for example, when the dynamical system is chaotic Brunton et al. (2016). Furthermore it may not be possible to observe the entirety of $X$. This may be handled as a missing data problem Little and Rubin (2019), and could be handled by joint parameter-state estimation for example using data assimilation; see Bocquet et al. (2019); Chen et al. (2022).
An additional complication with using direct data is ensuring the stability of the dynamical system with the calibrated error model $\delta(\cdot)$. Although with direct data, we can get more control of the accuracy of the error model itself, the calibrated error model often leads to unstable simulations of the dynamical system with the error model Beck et al. (2019); Brenowitz et al. (2020). There are several ways to mitigate the instability introduced by the error model, e.g., adopting a structure that ensures physical constraints Yuval et al. (2021), enforcing physical constraints Beucler et al. (2021), ensuring stability by bounding the eigenvalues of the linearized operator, and limiting the Lipschitz constant of $\delta(\cdot)$ Shi et al. (2019). However, a systematic approach to ensure stability is lacking.
@@ 2.1.2 Indirect Data
Instead of assuming access to direct data $\{X,\delta(X)\}$, the error model can also be calibrated with indirect data by solving an inverse problem associated with (6) (i.e., solve for the most likely parameters $\theta_{\textrm{P}},\theta_{\textrm{I}}$ given the model $\mathcal{G}$ and data $y$). Using indirect data involves simulating the dynamical system with the error model as in (5); therefore, the calibration procedure with indirect data favors error models that lead to stable simulations, an important advantage over the direct methods. Typical examples of problems giving rise to indirect data include time-series of $X$ for which the resolution is not fine enough to extract direct data for calibration Boys et al. (2008), time-averaged statistics of $X$ Schneider et al. (2017) or dynamical systems that are partially observed Pokern et al. (2009). More generally, indirect data can also be interpreted as constraints on $X$, and thus physical constraints can be enforced via augmenting indirect data.
@@ 2.2 Methods of Calibration
Using direct data $\{X,\delta(X)\}$ for calibration leads to a regression problem, which can be solved with standard methods for a given parameterization of the error model (e.g., least squares fit for dictionary learning, gradient descent methods for neural network). By contrast, using indirect data for calibration leads to an inverse problem associated with Eq. (6). Indirect methods can be linked to direct methods by framing the inverse problem as a missing data problem Little and Rubin (2019) and alternating between updating the missing data and updating the calibration parameters using learned direct data, for example, using the expectation-maximization (EM) algorithm Meng and Van Dyk (1997). However, in this section we focus on the calibration in the inverse problem setting; we discuss gradient-based and derivative-free optimization methods and how to enforce constraints.
@@ 2.2.1 Gradient-based or Derivative-free Optimization
is calculated, where $\theta=[\theta_{\textrm{I}},\theta_{\textrm{P}}]$ collects all parameters. In practice, the action of the gradient $d\mathcal{G}^{T}/d\theta$ is often evaluated via adjoint methods for efficiency. Although the gradient-based optimization is usually more efficient when $\mathcal{G}$ is differentiable, the evaluation of $\mathcal{G}$ can be noisy (e.g., when using finite-time averages to approximate infinite-time averaged data Dunbar et al. (2022)) or stochastic (e.g., when using stochastic processes to construct the error model). In these settings, gradient-based optimization may no longer be suitable, and derivative-free optimization becomes necessary. In this paper we focus on Kalman-based derivative-free optimization for solving the inverse problem; A briefly reviews a specific easily implementable form of ensemble Kalman inversion (EKI), to illustrate how the methodology works, and gives pointers to the broader literature in the field.

### v2

@@ 1 Introduction
Numerical simulation is at the heart of modeling, predicting, and understanding dynamical systems that are too complex to be amenable to analytical solution. Complex dynamical systems here extend from molecular dynamics with quantum effects to the planetary scales of weather and climate. The range of dynamically important scales in these systems can be vast, for example, in case of the atmosphere, extending over 13 orders of magnitude from the micrometers of cloud droplets and aerosols to the tens of thousands kilometers of planetary waves. The number of degrees of freedom that would need to be resolved for a faithful simulation of such systems (e.g., $\gtrsim 10^{21}$ for a typical atmospheric boundary layer flow) often exceeds what will be computationally feasible for the foreseeable future [1].
@@
where $f$ may depend on derivatives of the state $X$; hence, the system may represent partial differential equations. The system depends on empirical parameters $\theta_{\textrm{P}}$ that appear in closure models. For example, in large-eddy simulations of turbulent flows, the most energetic “large eddies” are explicitly resolved in the dynamics represented by $f$. The effect of the unresolved scales is modeled by subgrid-scale models, such as the classical Smagorinsky model [2], which depend on empirical parameters $\theta_{\textrm{P}}$ (e.g., the Smagorinsky coefficient). Similar semi-empirical models are used in many other fields. They encode domain-specific knowledge, and their parameters $\theta_{\textrm{P}}$ need to be calibrated with data.
@@
The challenge is that simulated observables $\hat{y}$ generally are biased estimates of actual data $y$. The actual data $y$ are affected by measurement error, and the simulated data $\hat{y}$ are affected by structural errors in the approximate dynamical system (1); both $y$ and $\hat{y}$ can also be affected by sampling error. For example, while a general feature of turbulence is to enhance mixing of conserved quantities, turbulent mixing is not always diffusive in character. Therefore, diffusive subgrid-scale models such as the Smagorinsky model are not always structurally correct, especially in convective situations with coherent flow structures [3]. This can lead to biases that, for example, adversely affect the calibration of model parameters $\theta_{\textrm{P}}$.
The purpose of this paper is to summarize principles of, and algorithms for, learning about structural error models that correct semi-empirical closure models. Wholesale replacement of semi-empirical closure models with neural networks and other deep learning approaches promises to overcome the structural strictures of existing closure models through more expressive models; it has recently received much attention [4, 5, 6, 7, 8, 9, 10, 11]. However, existing semi-empirical closure models encode valuable domain-specific knowledge. Learning flexible corrections to these models is often less data hungry, more interpretable, and potentially more generalizable than replacing them wholesale.
@@
Our approach to constructing error models builds upon but goes beyond the classical work of Kennedy and O’Hagan [12], who accounted for model error through an external bias correction term $\delta(X;\theta_{\textrm{E}})$, parameterized by parameters $\theta_{\textrm{E}}$ and inserted at the boundary between output from a computer model $\hat{y}=\mathcal{G}(\theta_{\textrm{P}})$ and data $y$:
@@
Here, $\mathcal{G}(\theta_{\textrm{P}})=\mathcal{H}[X(\theta_{\textrm{P}})]$ corresponds to solving (1) for the time series of the state $X$, which depends parametrically on $\theta_{\textrm{P}}$, and then applying the observation operator (2); hence, $\mathcal{G}$ is a mapping from the space of model parameters $\theta_{\textrm{P}}$ to the space of observations $y$. The noise $\eta$ represents additional (e.g., observation) errors, assumed to have zero mean. In this approach, the model parameters $\theta_{\textrm{P}}$ remain fixed (i.e., a property of $\mathcal{G}$) while parameters $\theta_{\textrm{E}}$ in the error model $\delta$ are tuned such that the residual $y-\hat{y}$ has a small magnitude and zero mean. This approach of externalizing model error for bias correction has been applied and further expanded in many subsequent papers [e.g., 13, 14, 15, 16, 17]. A key advantage of the external model error approach is that the model producing $\hat{y}$ can be treated as a black box, which facilitates use of this approach across different domains. State variables $X$ and residuals $y-\hat{y}$ form input-output pairs from which the error model $\delta(X;\theta_{\textrm{E}})$ can be learned, for example, with supervised learning [^3] Supervised learning refers to regression or interpolation of data. approaches, usually in an offline setting separate from learning about the model parameters $\theta_{\textrm{P}}$.
@@
    It is difficult to incorporate physical (or other process-based) knowledge or constraints (e.g., conservation laws) in the error model $\delta(X;\theta_{\textrm{E}})$ [15].
@@
To address these drawbacks, a few researchers have started to explore an approach that internalizes model error [18, 19, 20]. Such an internal model error approach embeds $L$ error models $\delta_{l}(X;\ \theta_{\textrm{I}}^{(l)})$ ($l=1,\dots,L$) within the dynamical system, at the places (e.g., in closure models) where the errors are actually made. Let their collection be written as
@@
Such an approach has found applications, for example, in turbulence modeling [21, 22, 23, 24, 25, 26, 27]. By incorporating the structural error models $\delta(X;\ \theta_{\textrm{I}})$ inside the dynamical system, the error models can in principle lead to improved predictions even of quantities that were not used to train the error models. The error models can be learned alongside the model parameters $\theta_{\textrm{P}}$ in an online setting. They also are more amenable to interpretation because they are included in the places where errors are actually made. A potential downside of internalizing model error is that the effects of the structural errors $\delta$ map onto data $y$ only through the resolved state $X$ of the dynamical system, and residuals $y-\hat{y}$ are generally not directly informative about how the structural errors $\delta(X;\ \theta_{\textrm{I}})$ depend on state variables $X$; thus, learning about structural errors $\delta(X;\ \theta_{\textrm{I}})$ can generally not be accomplished with direct supervised learning approaches. Instead, residuals $y-\hat{y}$ only provide indirect information about structural errors $\delta(X;\ \theta_{\textrm{I}})$. Additionally, if derivatives of the dynamical system $f$ with respect to parameters are not easily available, or if the dynamical system is not differentiable, gradient-based methods for learning about the model errors $\delta(X;\ \theta_{\textrm{I}})$ are difficult or impossible to use.
@@
    Gaussian processes, as in Kennedy and O’Hagan [12];
@@
    Models assembled from dictionaries of terms (e.g., involving differential operators), as in the data-driven discovery of partial differential equations [28, 29, 30, 31, 32, 33];
@@
    Neural networks, for their expressivity [4, 34, 6, 7, 35, 36];
@@
    Stochastic models, because without a clear scale separation between resolved and unresolved degrees of freedom, theoretical analysis suggests that closure models generally should be stochastic [e.g., 37, 38, 39, 40];
@@
    Non-local models, because structural errors may be non-local in space [41, 42, 43, 44, 45], in time [46, 47, 48], or in both [49].
We will discuss how to learn about such error models both from direct and indirect data. Supervised learning of various types of error models $\delta_{l}$ from direct data has been performed in a number of settings, for example, to discover terms in differential equations or neural network closure models from residual time tendencies that give direct information about the error model that is sought [e.g., 50, 26, 27, 51]. However, data directly informative about error models, such as high-resolution time tendencies, are not always available. When training an error model on time tendencies, it can also be difficult to ensure both stability of the dynamical system including the error model and satisfaction of physical constraints (e.g., energy conservation) [e.g., 36]. Training an error model with indirect data, in an inverse problem rather than supervised learning setting [e.g., 24, 52, 53], can be advantageous in those circumstances. We demonstrate how it can be accomplished.
@@
The principles and algorithms we will discuss are broad purpose and applicable across a range of domains and model complexities. To illustrate their properties in a relatively simple setting, we use two versions of the Lorenz 96 [54] dynamical system: the basic version of the model and its multiscale generalization [55]. Section 2 discusses the calibration of internal error models with direct or indirect data and the enforcement of constraints such as conservation properties. Section 3 introduces various ways of constructing error models. Section 4 introduces the two Lorenz 96 systems and then proceeds to present various concepts and methods through numerical results for these systems. The Lorenz 96 systems serve as illustrative examples of ergodic systems, for which we avoid trajectory matching and learn about unknown parameters from time-averaged statistics, which approximate the expectation with respect to the invariant measure. Section 5 is organized similarly to the previous section, but deals with non-ergodic systems for which trajectory matching is needed to learn about the underlying unknown parameters; as an example, we study a model of the human glucose-insulin system. In Section 6, we state our conclusions.
@@ 2.1.1 Direct Data
as an example. A fine temporal resolution of $X(t)$ is usually needed to approximate $\dot{X}-f(X;\theta_{\textrm{P}})$ and obtain estimates of the error $\delta(X;\theta_{\textrm{I}})$ as a residual. With this method, it becomes challenging to obtain reliable direct data when the trajectories $\dot{X}$ are noisy, for example, when the dynamical system is chaotic [50]. Furthermore it may not be possible to observe the entirety of $X$. This may be handled as a missing data problem [56], and could be handled by joint parameter-state estimation for example using data assimilation; see [57, 58].
An additional complication with using direct data is ensuring the stability of the dynamical system with the calibrated error model $\delta(\cdot)$. Although with direct data, we can get more control of the accuracy of the error model itself, the calibrated error model often leads to unstable simulations of the dynamical system with the error model [59, 60]. There are several ways to mitigate the instability introduced by the error model, e.g., adopting a structure that ensures physical constraints [61], enforcing physical constraints [62], ensuring stability by bounding the eigenvalues of the linearized operator, and limiting the Lipschitz constant of $\delta(\cdot)$ [63]. However, a systematic approach to ensure stability is lacking.
@@ 2.1.2 Indirect Data
Instead of assuming access to direct data $\{X,\delta(X)\}$, the error model can also be calibrated with indirect data by solving an inverse problem associated with (6) (i.e., solve for the most likely parameters $\theta_{\textrm{P}},\theta_{\textrm{I}}$ given the model $\mathcal{G}$ and data $y$). Using indirect data involves simulating the dynamical system with the error model as in (5); therefore, the calibration procedure with indirect data favors error models that lead to stable simulations, an important advantage over the direct methods. Typical examples of problems giving rise to indirect data include time-series of $X$ for which the resolution is not fine enough to extract direct data for calibration [64], time-averaged statistics of $X$ [65] or dynamical systems that are partially observed [66]. More generally, indirect data can also be interpreted as constraints on $X$, and thus physical constraints can be enforced via augmenting indirect data.
@@ 2.2 Methods of Calibration
Using direct data $\{X,\delta(X)\}$ for calibration leads to a regression problem, which can be solved with standard methods for a given parameterization of the error model (e.g., least squares fit for dictionary learning, gradient descent methods for neural network). By contrast, using indirect data for calibration leads to an inverse problem associated with Eq. (6). Indirect methods can be linked to direct methods by framing the inverse problem as a missing data problem [56] and alternating between updating the missing data and updating the calibration parameters using learned direct data, for example, using the expectation-maximization (EM) algorithm [67]. However, in this section we focus on the calibration in the inverse problem setting; we discuss gradient-based and derivative-free optimization methods and how to enforce constraints.
@@ 2.2.1 Gradient-based or Derivative-free Optimization
is calculated, where $\theta=[\theta_{\textrm{I}},\theta_{\textrm{P}}]$ collects all parameters. In practice, the action of the gradient $d\mathcal{G}^{T}/d\theta$ is often evaluated via adjoint methods for efficiency. Although the gradient-based optimization is usually more efficient when $\mathcal{G}$ is differentiable, the evaluation of $\mathcal{G}$ can be noisy (e.g., when using finite-time averages to approximate infinite-time averaged data [68]) or stochastic (e.g., when using stochastic processes to construct the error model). In these settings, gradient-based optimization may no longer be suitable, and derivative-free optimization becomes necessary. In this paper we focus on Kalman-based derivative-free optimization for solving the inverse problem; A briefly reviews a specific easily implementable form of ensemble Kalman inversion (EKI), to illustrate how the methodology works, and gives pointers to the broader literature in the field.
It is worth noting that EKI requires ensemble simulations, which may be less attractive when the system is not chaotic and gradient information is available. Even for chaotic systems, matching short trajectories can be achieved by gradient-based optimization [69], without the need to run ensemble simulations. For large datasets, the computational cost of EKI increases, e.g., because of the cost of the matrix inversions in the EKI formula. However, it is possible to make use of the sparse structure of the covariance matrix in data space and thus reduce the computational cost of the matrix inversion [70]. For a large dataset and a large number of parameters, a relatively small ensemble size may not suffice to approximate the covariance matrix; in that case, localization techniques [71, 72, 73] can be used to avoid spurious correlations due to the small ensemble size. In practice, localization techniques often lead to sparse covariance matrices, which allows users to employ standard efficient algorithms for the inverse and the multiplication of those covariance matrices in EKI.